OpenStudy (anonymous):

y = 3x - 4 y = -2x + 1

OpenStudy (anonymous):

do you know how to use substitution?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Solve the system of equations below by graphing them with a pencil and paper. Enter your answer as an ordered pair. y = 3x - 4 y = -2x + 1

OpenStudy (anonymous):

oh you have to solve by graphing. okay so the last number represents the y-intercept and the 3x and -2x is the slope

OpenStudy (amorfide):

let y=3x-4 be equation 1 let y=-2x+1 equation 1 is equal to y equation 2 is equal to y therefore we know that equation 1 equals equation 2 3x-4=2x+1

OpenStudy (amorfide):

oh you gotta graph lol

OpenStudy (anonymous):